OpenStudy (anonymous):

Question on groups and symmetries Let G = {(x, y, z): x, y, z are elements on R } and define (x1, y1, z1)*(x2, y2, z2) = (x1+x2, y1+x1z2+y2, z1+z2) 1) Show that (G,*) is a group and that it is non-abelian 2) Show that (G,*) does not have any elements of finite order (other than the identity)

OpenStudy (anonymous):

@satellite73 for this question it is telling us to show if its a group or not now we have to show if associativity, identitym and the inverse hold?

OpenStudy (anonymous):

yeah have fun

OpenStudy (anonymous):

alright you know how it says (G,*) i dont know what that * defines ?

OpenStudy (anonymous):

you wrote what it defines

OpenStudy (anonymous):

look carefully at the question you posted. it says right in it what "*" is

OpenStudy (anonymous):

(x1, y1, z1)*(x2, y2, z2) = (x1+x2, y1+x1z2+y2, z1+z2) ?

OpenStudy (anonymous):

that is what you wrote right?

OpenStudy (anonymous):

yeahh

OpenStudy (anonymous):

then that is what * is

OpenStudy (anonymous):

associative is going to be totally straight forward but a real pain to write. you have to show \[\left((x_1,y_1,z_1)*(x_2,y_2,z_2)\right)*(x_3,y_3,z_3)=(x_1,y_1,z_1)*\left((x_2,y_2,z_2)*(x_3,y_3,z_3)\right)\] it is going to be a lot of writing, but not much else

OpenStudy (anonymous):

get a nice sharp pencil

OpenStudy (anonymous):

lol and that is just (x1,y1,z1)*(x2+x3, y2+x2z3+y3, z2+z3)

OpenStudy (anonymous):

and now we multiply this guy again to get (x1+x2+x3, y1+x1z2+x1z3+y2+x2z3+y3, z1+z2+z3)

OpenStudy (anonymous):

is that all we do for associativity ?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

you have to do it twice to show that they are the same

OpenStudy (anonymous):

you did \((ab)c\) now you have to do \(a(bc)\) and show that they are equal no doubt they will be

OpenStudy (anonymous):

oooo ok ok yeah your right

OpenStudy (anonymous):

oh now that i look closer it seems you did \(a(bc)\) so now you have to do \((ab)c\) i said it is a pain

OpenStudy (anonymous):

loll

OpenStudy (anonymous):

you just have to write it. there is nothing really else to do

OpenStudy (anonymous):

yeah truee next to find the identity we this is what throws me off for identity ae=a and ea=a is our identity just zero?

OpenStudy (anonymous):

sorry would it be 1 because anything multiplied to itself has to be that same specific number?

OpenStudy (anonymous):

@Zarkon could you pleaseeeee help me on this questionn

OpenStudy (anonymous):

i really need help on this final questionn

OpenStudy (anonymous):

@Zarkon @TuringTest pleasee help

OpenStudy (anonymous):

@satellite73 could you please look at this question and help me on this proof pleaseeee